• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Zheng

    Log entry time 15:50:44 on March 3,2009

    Entry number 264340

    keyword=Rebooted ROCs 5, 8, 9, 10 etc, and changed LHRS p0 to 1.13 (#F)

    To fix the BB sync problem we rebooted a bunch of ROCs. THis happened after run 1951. When rebooting we probablly also accidentally rebooted some of the LHRS TS or ROCs so we killed LHRS CODA too. This happened at the end of run 20498 (hence no end of run entry).

    WHile Bob is fixing the BB coda, we changed the left HRS momentum to 1.13 GeV/c (Kine #F).

    15:30, both DAQs are up and we started BB run 1952.

    15:35, Left VDC HV tripped, Did Crate reset (VDC high voltage line A and B).

    15:40. Started LHRS run 20499 on Kinematics F.